OpenStudy (anonymous):

3 Times the square root of 45 ?

OpenStudy (unklerhaukus):

\[3 \times √45\] \[3 \times \sqrt{5 \times 9}\] \[3 \times \sqrt{5} \times 3\] \[9√5\]

OpenStudy (mertsj):